### v4

# Analytics, have some humility: a statistical view of fourth-down decision making

Abstract
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from historical data. These models, however, are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Thus, it is imperative to knit uncertainty quantification into the fourth-down decision procedure; we do so using bootstrapping. We find that uncertainty in the estimated optimal fourth-down decision is far greater than that currently expressed by sports analysts in popular sports media. Our contribution is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim.
Keywords: Decision making under uncertainty, applications and case studies, machine learning, bootstrap/resampling, statistics in sports
1 Introduction
In-game strategic decision making is one of the fundamental objectives of sports analytics. To mathematically compare strategies, analysts need a value function that measures the value of each game-state. The optimal decision maximizes the value of the next game-state. Across sports, however, value functions are not observable quantities; they are defined by models. It is the sports analysts’ task to infer the value of each game-state from the massive dataset of all plays in the recent history of a given sport.
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, 2009b suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke, 2009a suggest making the decision that maximizes estimated ${\mathsf{WP}}$. These analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
The ${\mathsf{WP}}$ estimates used for fourth-down decision making typically arise from statistical models fit from historical data. Given the results of the entire recent play-by-play history of football, these models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. However, these models are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. Because every game has only one winner, the binary win/loss response values are not independent, as all plays from the same game share the exact same draw of the win/loss outcome. This reduces the effective sample size and inflates standard errors.
Traditional decision procedures, based solely on effect size, ignore the uncertainty inherent in estimating ${\mathsf{WP}}$ from noisy and highly correlated historical data. This uncertainty should percolate into the fourth-down decision procedure, but popular approaches currently treat win probabilities as known quantities rather than estimates. Trusting the resulting ${\mathsf{WP}}$ point estimates to dictate fourth-down decision making without accounting for uncertainty in these estimates leads to overconfident and potentially sub-optimal decisions.
Our focus is not to “fix” win probability models by adjusting for additional covariates or reducing model bias. Rather, we shed light on the high variance nature of estimating win probability and show that such estimates are rife with uncertainty. We use bootstrapping to quantify uncertainty in fourth-down recommendations and recommend a fourth-down decision when we are confident it has higher ${\mathsf{WP}}$ than all other decisions. Our contribution to the literature is a major advance in fourth-down strategic decision making: far fewer fourth-down decisions are as obvious as analysts claim. A substantial proportion of fourth-down decisions feature ${\mathsf{WP}}$ recommendations subject to large uncertainty. Thus, we ask football analysts to have some humility: for many game-states, there is simply not enough data to trust ${\mathsf{WP}}$ point estimates.
The remainder of this paper is organized as follows. In Section 2 we detail the traditional fourth-down decision procedure. Then in Section 3 we incorporate uncertainty quantification into the fourth-down decision procedure. In Section 4 we present our main findings and we conclude in Section 5.
2 The traditional fourth-down decision procedure
The standard mathematical approach to fourth-down decision making is to make the decision that maximizes estimated win probability. In this Section we explore the traditional decision procedure and detail how analysts commonly estimate win probability today. We begin with a brief overview of our historical dataset of football plays in Section 2.1. Then in Section 2.2 we illustrate the traditional decision process for example plays. In Section 2.3 we describe how analysts estimate a team’s fourth-down win probability if it punts, kicks a field goal, or attempts a conversion. Fourth-down win probability is typically estimated in terms of first-down win probability, so in Section 2.4 we describe well known open source first-down win probability models.
2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that are relevant to estimating win probability that describe the context of the play, such as the score differential, time remaining, yardline, down, yards to go, etc. (see Table 3 in Appendix A for descriptions of relevant variables). Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety.
In order to avoid data bleed in constructing the various decision transition models (Appendix C) relevant for fourth-down decision making, we hold out plays from 1999 to 2005. Additionally, we hold out plays from 2022 to illustrate our decision procedure on out-of-sample example plays in Section 4.1. We are then left with a primary dataset of $600,825$ plays, with $240,380$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.111 https://github.com/snoopryan123/fourth_down
2.2 Example plays
The traditional fourth down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. Further, the strength of a decision is proportional to the estimated gain in win probability by making that decision. Notably, Baldwin, (2021) and Burke, 2009a employ this decision procedure and post their fourth-down recommendations on Twitter.222 Baldwin’s posts his fourth-down recommendations at @ben_bot_baldwin and Burke posts his at @bburkeESPN. We illustrate this procedure through two example plays.
Example play 1. First, we consider an example of Baldwin’s fourth-down decision charts, which is based on a model described in Section 2.4. Figure 1 illustrates Baldwin’s decision-making charts for a fourth-down play in which the Colts had the ball against the Patriots in Week 10 of 2023. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
Figure 1: Baldwin’s decision charts for example play 1. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal.
Example play 2. Next, we give an example of Burke’s decision chart, which is based on a proprietary model. Figure 2 illustrates Burke’s fourth-down decision chart for a play from the 2023 NFC Championship game. The chart visualizes the estimated optimal decision (color) according to effect size as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
Figure 2: Burke’s decision boundary chart for example play 2. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary.
2.3 Estimating the win probability of a fourth-down decision
An offense has three possible decisions to make on fourth down: it can punt the ball, kick a field goal, or attempt a conversion (colloquially known as “going for it”), denoted ${\mathsf{Punt}}$, ${\mathsf{FG}}$, and ${\mathsf{Go}}$, respectively. To employ the traditional fourth-down decision procedure, we need to estimate a team’s fourth-down win probability for each decision as a function of the game-state.
In Appendix B we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We can do this because a first down always follows a fourth down. For instance, success on a conversion attempt results in a first down for the team with possession and failure results in a first down for the other team. Thus, ${\mathsf{Go}}$ win probability on fourth down is a function of conversion success probability and first-down win probability. Similarly, ${\mathsf{FG}}$ win probability is a function of field goal success probability and first-down win probability. ${\mathsf{Punt}}$ win probability is a function of punt outcome and and first-down win probability. We estimate these decision transition probabilities in Appendix C.
We detail first-down win probability estimation in the main text, relegating the remainder of fourth-down win probability modeling to the Appendix, because it is the most difficult component of fourth-down decision making. Whereas decision transition models are fit from i.i.d. play-level outcomes and so have a relatively large effective sample size, first-down win probability models are fit from noisy, clustered, and highly correlated game-level outcomes. There are just $4,101$ games in our dataset, providing a relatively small effective sample size for statistical first-down win probability models. On this view, first-down win probability forms the crux of the fourth-down decision framework.
2.4 Estimating the win probability of a first down
Win probability estimates arise broadly from one of two classes of models, probabilistic state-space models or statistical models. On one hand, state-space models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and are then propagated into win probability by simulating games. When implemented correctly, these models are sensible ways to estimate ${\mathsf{WP}}$. However, they are difficult in practice, as they require: a careful encoding of the convoluted rules of football into a set of states and the actions between those states, careful estimation of transition probabilities, and enough computing power to run enough simulated games to achieve desired granularity. Each of these are nontrivial.
On the other hand, statistical models are fit entirely from historical data. Given the results of a set of observed football plays, statistical models fit the relationship between certain game-state variables using data-driven regression or machine learning approaches. These models are widely used today in the football analytics community because of publicly available play-by-play data (e.g., nflFastR (Carl and Baldwin,, 2022)) and accessible off-the-shelf machine learning models (e.g., ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016)). Additionally, due to a perceived abundance of data, flexible machine learning models are viewed as more “trustworthy” than previous mathematical models that make more simplifying assumptions. For these reasons, the open source win probability models used today for fourth-down decision making are statistical / machine learning models. Hence, in this paper we focus on statistical / machine learning win probability models.
Lock and Nettleton, (2014) use a Random Forest (${\mathsf{RF}}$) (Breiman,, 2001) to estimate win probability from historical data. The response variable is a binary variable indicating whether the team with possession wins the game. They model win probability as a function of score differential, game seconds remaining, yardline, down, yards to go, the number of timeouts remaining for each team, pre-game point spread, total points scored, and an additional feature to capture the change in impact of score differential over the course of a game,
adjusted score=score differential1+game seconds $\text{adjusted score}=\frac{\text{score differential}}{\sqrt{1+\text{game seconds remaining}}}.$ (1)
They use a Random Forest of 500 regression trees with parameters ${\mathsf{mtry}}=2$ and ${\mathsf{nodesize}}=200$.
Baldwin, (2021) uses ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016) to estimate win probability from historical data. He uses the same binary win/loss response variable as before. Baldwin models win probability as a function of score differential, game seconds remaining, half seconds remaining, yardline, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, and the number of timeouts remaining for each team. He uses two additional features to capture the change in impact of point spread and score differential over the course of a game,
[…]
and
[…]
Baldwin includes monotonic constraints for yardline, yards to go, down, score differential, timeouts remaining for each team, spread-time, and diff-time-ratio. He tunes the ${\mathsf{XGBoost}}$ hyper-parameters by minimizing cross validated log-loss (Baldwin,, 2021).
We also initially considered the generalized additive model (${\mathsf{GAM}}$) (Hastie and Tibshirani,, 1986) win probability model from Yurko et al., (2018) but opted in this paper to focus just on flexible non-parametric models because we expect win probability to be replete with interactions between variables.
Each of these win probability models are statistical models fit from an historical play-by-play dataset. This dataset includes plays from all downs; these models fit first-down win probability by including down as a covariate. We find that a model fit from just first-down plays better predicts first-down win/loss (see Appendix D), which is what we care about for fourth-down decision making. For the decision procedure used in the remainder of this paper, we use a first-down ${\mathsf{XGBoost}}$ win probability model fit from just first downs as described in Appendix D.
3 Fourth-down decision making under uncertainty
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. The win probability models that form the crux of traditional fourth-down decision procedures today are machine learning models fit from historical data. Football analysts today typically use these models because they see a massive dataset – $600,825$ plays with $240,380$ first-down plays in our play-by-play dataset – and think this is enough data to fit a highly accurate model using machine learning. However, this is not true because the binary win/loss response variable is noisy and highly correlated: every game has only one winner. Formally, the binary response variable $y_{i}$ of the $i^{th}$ play indicates whether the team with possession won the game. The response values are not independent, as all plays from the same game share the same draw of the response column. Thus the effective sample size is somewhere between the number of plays ($600,825$) and the number of non-tied games ($4,101$) in our dataset. This is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context, and as such we expect win probability point estimates to have high variance.
3.1 Simulation study: random walk football
To illustrate just how difficult it is to accurately fit a statistical win probability model from noisy and highly correlated observational data, we conduct a simulation study. We create a simplified random walk version of football in which the true win probability at each game-state is known. Then, we see how well a statistical ${\mathsf{WP}}$ model recovers the true win probability. These models find the general ${\mathsf{WP}}$ trend (i.e., they they have low bias) but are subject to substantial uncertainty: to achieve adequate frequentist marginal coverage, bootstrapped ${\mathsf{WP}}$ confidence intervals need to be substantially wide. As real football is exponentially more complex, its confidence intervals should be even wider.
Rules of random walk football. Random walk football begins at midfield. Each play, the ball moves left or right by one yardline with equal probability. If the ball reaches the left (right) end of the field, team one (two) scores a touchdown, worth $+1$ ($-1$) point. The ball resets to midfield after each touchdown. After $N$ plays, the game ends. If the game is still tied after $N$ plays, a fair coin is flipped to determine the winner. We include the formal mathematical specification of the game in Appendix E. We also explicitly compute true win probability as a function of time, field position, and score differential using dynamic programming in Appendix E.
Simulation methodology. One simulated simplified football dataset consists of $G$ games, each with $N$ plays per game. We use $L=4$ yardlines so that the average number of first-down plays between each score is similar to that of a real football game. $M=25$ times we simulate a training dataset ${\mathscr{D}}_{\text{train}}$ and a testing dataset ${\mathscr{D}}_{\text{train}}$, each of which is a draw of simplified football history. Each of the $M$ simulated training and testing datasets has the form
${\mathscr{D}}_{\text{train}}=\{(n,X_{gn},S_{gn},y_{gn}):n=1,...,N\text{ and }g=1,...,G\}.$ (4)
and
${\mathscr{D}}_{\text{test}}=\{(n,X_{gn},S_{gn},{\mathsf{WP}}_{gn}):n=1,...,N\text{ and }g=1,...,G\},$ (5)
respectively. For each play of game $g$ we record the timestep $n$, the field position $X_{gn}$, the score differential $S_{gn}$, a binary variable $y_{gn}$ indicating whether the team with possession wins the game, and the true win probability ${\mathsf{WP}}_{gn}$. The response variable $y$ is clustered and highly correlated, as each play within the same game shares the same random draw of $y$. From the training set we fit a machine learning model to estimate win probability as a function of timestep $n$, field position $x$, and score differential $s$,
$\widehat{{\mathsf{WP}}}(n,x,s)={\mathsf{XGBoost}}({\mathscr{D}}_{\text{train}})(y|n,x,s).$ (6)
We use half of the games from the training set as a validation set to tune the ${\mathsf{XGBoost}}$ model. We also create various bootstrapped win probability confidence intervals, discussed below, denoted $CI(n,x,s)$.
For each simulation we compute the mean absolute error between true and estimated win probability on the test set,
${\mathsf{MAE}}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}|\widehat{{\mathsf{WP}}}(n,x_{gn},s_{gn})-{\mathsf{WP}}_{gn}|.$ (7)
Also, for each simulation and each confidence interval method, we compute the pointwise marginal coverage
$\text{coverage}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}\mathbb{I}\left({\mathsf{WP}}_{gn}\in CI(n,x_{gn},s_{gn})\right)$ (8)
and mean length
mean $\text{mean length}=\frac{1}{G\cdot N}\sum_{(n,g)\in{\mathscr{D}}_{\text{test}}}|CI(n,x_{gn},s_{gn})|.$ (9)
We report the ${\mathsf{MAE}}$, CI marginal coverage, and CI mean length averaged over the $M$ simulations and include plus/minus twice their standard errors.
Bootstrap confidence interval methodology. We compare the coverage and lengths of ${\mathsf{WP}}$ confidence intervals produced by the standard bootstrap, cluster bootstrap, and randomized cluster bootstrap, averaged over the $M$ simulations. In the standard bootstrap, which assumes each row (play) of the dataset is independently drawn, each of $B$ bootstrapped datasets are formed by re-sampling $N$ plays uniformly with replacement. In the cluster bootstrap, each of $B$ bootstrapped datasets are formed by re-sampling $G$ games uniformly with replacement, keeping each observed play within each re-sampled game. Finally, in the randomized cluster bootstrap, each of $B$ bootstrapped datasets are formed by re-sampling $G$ games uniformly with replacement, and within each game re-sampling $N$ plays uniformly with replacement.
Then, for each bootstrap method, we fit a ${\mathsf{WP}}$ model ${\mathsf{WP}}_{b}$ to each bootstrapped dataset $b=1,...,B$. The $90\%$ confidence interval for the ${\mathsf{WP}}$ estimate at game-state ${\mathbf{x}}$ is defined by the $5^{th}$ and $95^{th}$ quantiles of ${\mathsf{WP}}_{1}({\mathbf{x}}),...,{\mathsf{WP}}_{B}({\mathbf{x}})$. We use $B=100$ in this simulation study. Also, to avoid substantially low coverage near the extremes (${\mathsf{WP}}\approx 0$ or ${\mathsf{WP}}\approx 1$), we widen our confidence intervals when $\widehat{{\mathsf{WP}}}<0.025$ to have a lower bound of 0 and when $\widehat{{\mathsf{WP}}}>0.975$ to have an upper bound of 1.
Initial simulation results. In Table 1 we report the initial results of our simulation study. In the second column of Table 1, each simulated dataset consists of $G=4,101$ games and $N=53$ plays per game, which matches the number of games and the average number of first-down plays in our dataset of real football plays. Each game in each of these datasets consists of $K=53$ plays per game that share the same outcome. In the third column of Table 1, each simulated dataset consists of $G=4,101\cdot 53$ games and $N=53$ plays per game. Then, we remove all but $K=1$ play per game so that each timestep $n$ has exactly $4,101$ corresponding i.i.d. rows in the dataset. In other words, those datasets consist of $217,353$ plays with i.i.d. outcomes. For the simulation studies in each column, we report the mean absolute error of win probability point estimates. We also report the marginal coverage and mean length of $90\%$ confidence intervals produced by our three bootstrap methods. Alongside each value, which is an average over the $M$ simulations, we report plus/minus twice its standard error.
$G=4,101$,  $N=53$,  $K=53$ $G=4,101\cdot 53$,  $N=53$,  $K=1$
MAE bt ${\mathsf{WP}}$ and $\widehat{\mathsf{WP}}$ $0.0159\pm 0.0005$ $0.0155\pm 0.0004$
$90\%$ SB CI coverage $0.60\pm 0.029$ $0.65\pm 0.011$
$90\%$ CB CI coverage $0.74\pm 0.029$
$90\%$ RCB CI coverage $0.80\pm 0.028$
$90\%$ SB CI mean length $0.028\pm 0.001$ $0.029\pm 0.001$
$90\%$ CB CI mean length $0.047\pm 0.001$
$90\%$ RCB CI mean length $0.056\pm 0.001$
Table 1: Initial simulation study results for $G=4,101$ $N=53$, and $K=53$ (second column) and $G=4,101\cdot 53$ $N=53$, and $K=1$ (third column). We report the ${\mathsf{MAE}}$ (Formula (7)) of ${\mathsf{XGBoost}}$ ${\mathsf{WP}}$ point estimates and the marginal coverage (Formula (8)) and mean length (Formula (9)) of $90\%$ ${\mathsf{WP}}$ confidence intervals from the standard bootstrap (SB), cluster bootstrap (CB), and randomized cluster bootstrap (RCB). We include these values averaged over the $M$ simulations plus/minus twice their standard errors.
In the simulation study with highly correlated outcomes ($K=53$), the mean absolute error (${\mathsf{MAE}}$) between the true and estimated ${\mathsf{WP}}$ is less than $2\%$ over average. This indicates ${\mathsf{XGBoost}}$ recovers the general trend of true ${\mathsf{WP}}$, which we visualize in Figure 3a. In the simulation study with independent observations ($K=1$), the ${\mathsf{MAE}}$ is smaller. This suggests that some bias is induced by fitting ${\mathsf{WP}}$ from highly correlated observational data. Nonetheless, the ${\mathsf{MAE}}$ between the highly correlated ($K=53$) and independent ($K=1$) simulation studies are overall quite similar. This suggests that most of the error accrued when estimating ${\mathsf{WP}}$ from observational data is due to having limited data and a noisy binary response column.
Figure 3: Win probability ($y$-axis) as a function of play number $n$ ($x$-axis) for various values of score differential $s$ (color) at yardline $x=3$. True win probability is the solid line. On the left: for one of the simulations, ${\mathsf{WP}}$ estimates (dotted line) get the general trend right. On the right: $90\%$ bootstrapped ${\mathsf{WP}}$ confidence intervals (shaded regions) need to be substantially wide to achieve $90\%$ marginal frequentist coverage.
The marginal frequentist coverage of ${\mathsf{WP}}$ confidence intervals is also impacted by the correlated nature of football play-by-play data. The standard bootstrap achieves $5\%$ higher coverage for $K=1$ than for $K=53$. Since rows are independent in the simulation study without auto-correlation, each bootstrap method is identical when $K=1$. In the simulation study with highly correlated outcomes ($K=53$), mimicking observed football data, $90\%$ win probability confidence intervals yield much less than $90\%$ coverage. The best marginal coverage is a subpar $80\%$ from the randomized cluster bootstrap, whose $90\%$ confidence intervals have a large mean length $5.6\%$ ${\mathsf{WP}}$.
Tuning the bootstrap. Even in the obscenely simplified and idealized setting of random walk football, typical bootstrapped win probability confidence intervals don’t achieve proper coverage. The naive standard bootstrap in particular achieves dismally low coverage. Even the randomized cluster bootstrap that accounts for the correlation structure of observational football data doesn’t achieve high enough coverage. This coverage issue is likely worse for real football data, which is exponentially more complex than random walk football.
Since naive bootstrapped confidence intervals are not wide enough, we wish to tune the bootstrap so that its $90\%$ confidence intervals indeed produce $90\%$ marginal coverage. The traditional method of tuning non-parametric bootstrapped confidence intervals is to calibrate the bootstrapped quantiles (DiCiccio and Efron,, 1996). For instance, instead of using the $\alpha/2^{th}$ and $(1-\alpha/2)^{th}$ quantiles of ${\mathsf{WP}}_{1}({\mathbf{x}}),...,{\mathsf{WP}}_{B}({\mathbf{x}})$ to form a $1-\alpha$ confidence interval, we could use the $\beta/2^{th}$ and $(1-\beta/2)^{th}$ quantiles for some $\beta<\alpha$.
Although calibrating the bootstrapped quantiles fixes the coverage issue for ${\mathsf{WP}}$ confidence intervals, it doesn’t calibrate the bootstrap percentage, a metric that we introduce later in Section 3.2 to measure uncertainty in the estimated optimal fourth-down decision. Therefore, we use an alternate method to tune the bootstrap that calibrates both ${\mathsf{WP}}$ confidence intervals and the bootstrap percentage. Specifically, we introduce a tuning parameter $\phi\in(0,1]$ representing the fraction of data to be re-sampled in generating a bootstrapped dataset. By re-sampling less data than in the original dataset, we widen bootstrapped confidence intervals and increase marginal coverage. In the $\phi$-standard bootstrap, we re-sample $N\cdot\phi$ plays (rows) uniformly with replacement. In the $\phi$-cluster bootstrap, we re-sample $G\cdot\phi$ games uniformly with replacement, keeping each observed play within each re-sampled game. Finally, in the $\phi$-randomized cluster bootstrap, we re-sample $G\cdot\phi$ games uniformly with replacement, and within each game re-sample $N$ plays uniformly with replacement.
$\phi$-simulation results. In Table 2 we report the results of our simulation study in which we use $\phi$ to tune our bootstraps. Each simulated dataset consists of $G=4,101$ games, $N=53$ plays per game, and $K=53$ plays per game that share the same outcome. We report the marginal coverage and mean length of $90\%$ confidence intervals produced by our three bootstrap methods with $\phi=1$ (second column) and $\phi=0.5$ (third column). Alongside each value, which is an average over the $M$ simulations, we report plus/minus twice its standard error. As expected, lowering $\phi$ widens the confidence intervals and increases marginal coverage. As desired, $90\%$ confidence intervals from the randomized cluster bootstrap with $\phi=0.5$ achieve $90\%$ marginal coverage. To achieve satisfactory marginal coverage, however, ${\mathsf{WP}}$ confidence intervals need to be substantially wide, which we visualize in Figure 3b. Specifically, the randomized cluster bootstrap with $\phi=0.5$ produces $90\%$ confidence intervals with a mean width of $7.4\%$ ${\mathsf{WP}}$. This result is striking: in the extremely simplified setting of random walk football, win probability confidence intervals need to be extremely wide to achieve adequate coverage. So, it is very difficult to accurately estimate win probability by fitting a machine learning model from noisy and highly correlated football game outcomes. These are high variance estimators subject to large uncertainty.
$\phi=1$ $\phi=0.5$
$90\%$ $\phi$-SB CI coverage $0.60\pm 0.029$ $0.69\pm 0.027$
$90\%$ $\phi$-CB CI coverage $0.74\pm 0.029$ $0.85\pm 0.027$
$90\%$ $\phi$-RCB CI coverage $0.80\pm 0.028$ ${\color[rgb]{1,0,1}\textbf{0.90}}\pm 0.026$
$90\%$ $\phi$-SB CI mean length $0.028\pm 0.001$ $0.036\pm 0.001$
$90\%$ $\phi$-CB CI mean length $0.047\pm 0.001$ $0.063\pm 0.001$
$90\%$ $\phi$-RCB CI mean length $0.056\pm 0.001$ ${\color[rgb]{1,0,1}\textbf{0.074}}\pm 0.001$
Table 2: Simulation study results for $\phi=1$ (second column) and $\phi=0.5$ (third column) with $G=4,101$, $N=53$, and $K=53$ for both. We report the marginal coverage (Formula (8)) and mean length (Formula (9)) of $90\%$ ${\mathsf{WP}}$ confidence intervals from the standard bootstrap (SB), cluster bootstrap (CB), and randomized cluster bootstrap (RCB). We include these values averaged over the $M$ simulations plus/minus twice their standard errors.
3.2 Quantifying uncertainty in the estimated optimal decision
The fourth-down decision procedure widely used today in the football analytics community is to make the decision that maximizes estimated win probability. Win probability estimates arise from machine learning models fit from observational football data, which consists of noisy and highly correlated game outcomes. As discussed in the previous Section 3.1, even in simplified random walk football, win probability models are high variance estimators subject to substantial uncertainty. ${\mathsf{WP}}$ models fit from real football data are likely subject to even more uncertainty since real football is so much more complex. Therefore, it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure.
Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty. The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. We specify the full technical details of how we generate each bootstrapped dataset in Appendix F and how we choose an appropriate value for $B$ in Appendix G. Each bootstrapped dataset mimics a new draw of the training set: keeping all of the pre-game attributes of each game fixed (e.g., each team’s composition), bootstrapping mimics re-drawing plays and the random win/loss outcome of each game. Concisely, each bootstrapped dataset represents a re-draw of the recent history of football. Then, to each bootstrapped dataset we fit a fourth-down win probability model (described in Section 2.3 and Appendix B). At each game-state, each model yields an estimated optimal decision. The strength of each estimated optimal decision is proportional to the effect size, or the estimated gain in win probability by making that decision. Because the plays and win/loss outcomes in each bootstrapped dataset differ from those in the original dataset, each bootstrapped decision function differs from the original one.
Given the original effect size $\hat{g}({\mathbf{x}})$ and the $B$ bootstrapped effect sizes $\hat{g}_{1}({\mathbf{x}}),...,\hat{g}_{B}({\mathbf{x}})$ at game-state ${\mathbf{x}}$, we construct a confidence interval on the effect size using the quantiles. For instance, if $B=100$, a $90\%$ confidence interval on the effect size is $[\hat{g}_{5}({\mathbf{x}}),\ \hat{g}_{95}({\mathbf{x}})]$. This confidence interval represents the following. Assume that the win/loss outcomes across football history are generated by some underlying (and unknown) win probability function. If we re-simulated the history of football say $100$ times from that underlying ${\mathsf{WP}}$ function, keeping the pre-game conditions of each game the same, the true win probability is expected to lie in the $90\%$ confidence interval $90$ times.
Since a confidence interval is defined by two numbers, it is difficult to meld those two numbers into one decision recommendation. For example, consider a play in which the estimated gain in win probability for ${\mathsf{Go}}$ is $1\%$. If our confidence interval for this gain is $[-3\%,5\%]$, it is not clear how strongly we should recommend ${\mathsf{Go}}$. Similarly, it is difficult to compare two confidence intervals of varying lengths. For example, it is not clear whether our recommendation for a confidence interval of $[-3\%,5\%]$ should be weaker than that for a confidence interval of $[-1\%,3\%]$. Further, our primary interest is to quantify uncertainty in the decision itself (Friedman et al.,, 1999), which is not as granular as quantifying uncertainty in the effect size.
To boil a bootstrapped confidence interval on the effect size into a single number to drive decision making, we bag the decision itself. Specifically, we consider the bootstrap percentage (${\mathsf{boot\%}}$), or the percentage of bootstrapped models that report each decision to be optimal. Formally, at game-state ${\mathbf{x}}$ we have the original estimated optimal decision $d({\mathbf{x}})\in\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ (the point estimate) and $B$ bootstrapped estimated optimal decisions $d_{1}({\mathbf{x}}),...,d_{B}({\mathbf{x}})$. Then
${\mathsf{boot\%}}({\mathbf{x}})=100\%\cdot\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\left(d_{b}({\mathbf{x}})=d({\mathbf{x}})\right).$ (10)
Bootstrap percentage quantifies how confident we are that a fourth-down recommendation has higher win probability than all the other decisions. A high bootstrap percentage (${\mathsf{boot\%}}\approx 100\%$) reflects high confidence in the estimated optimal decision: the estimated optimal decision remains the same across the vast majority of re-draws of the training set. When ${\mathsf{boot\%}}$ is high, we can trust the decision recommended by the point estimate. Conversely, a low bootstrap percentage reflects low confidence in the estimated optimal decision: the estimated optimal decision differs substantially across different re-draws of the training set. A low bootstrap percentage is ${\mathsf{boot\%}}\approx 50\%$ or lower because a fourth-down decision almost always boils down to picking between ${\mathsf{Go}}$ and one of the kicks, as it is clearly sub-optimal to punt near the opponent’s endzone and to attempt a field goal far from the opponent’s endzone. Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncrasies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal. In a “parallel universe” in which we re-drew each random win/loss game outcome, the estimated optimal decision at that game-state may differ solely due to the randomness inherent in drawing the history of football.
On this view, bootstrap percentage is a measure of data reliability. At each game-state the model produces a point estimate of the effectiveness of a fourth-down decision; bootstrap percentage tells us how reliable this estimate is, or how much the data trusts its own estimate. To understand, think of the outcome (winning team) of each row (play) in the dataset as a random draw. If some of these draws resulted in different outcomes, our fitted win probability functions would be different. The less data we have access to, the more sensitive models are to the random idiosyncrasies of any particular training dataset. The bootstrap quantifies this sensitivity: given the amount of data we have, it quantifies the spectrum of variability in potential resulting fitted models based only on sampling uncertainty, or uncertainty resulting from having limited data.
4 Results
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure ignores the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Thus, we modified the decision procedure to account for uncertainty: we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Specifically, we measure our confidence in a decision by the proportion of bootstrapped win probability models that report the decision as optimal.
4.1 Example plays
We illustrate our decision procedure using example plays. Also, to compare our decision procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We detail the specification of this baseline coach model in Appendix H and include the model’s predictions in our decision figures.
Example play 1. First, we compare Baldwin’s fourth-down decision making procedure to ours via example play 1 from Section 2.2. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal, shown in Figure 1. In Figure 4 we incorporate uncertainty quantification into decision making. Although our point estimate (the blue column) suggests that ${\mathsf{Go}}$ provides a $1.4\%$ gain in win probability over ${\mathsf{FG}}$, our $90\%$ confidence interval $[-3.9\%,5.1\%]$ suggests that ${\mathsf{Go}}$ could either be a terrible or a great decision. Also, $50\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column), reflecting massive uncertainty in the optimal fourth-down decision. So, we are not confident in the fourth-down recommendation provided by the point estimate.
Figure 4: Our decision chart for example play 1. The bootstrapped models are split amongst ${\mathsf{Go}}$ and ${\mathsf{FG}}$ (${\mathsf{boot\%}}$, in the orange column, is $50\%$-$50\%$), so we don’t have enough data to be confident in the fourth-down recommendation provided by the point estimate (blue column).
Example play 2. Next, we compare Burke’s fourth-down decision procedure to ours via example play 2 from Section 2.2. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot in Figure 2 (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary. That chart, however, says nothing about the estimated strength of making the optimal decision or about uncertainty quantification. Hence, in Figure 5a we show our version of Burke’s chart, which uses color intensity to visualize the estimated gain in win probability by making a decision (darker colors indicate larger values).333 We use green for ${\mathsf{Go}}$, yellow for ${\mathsf{FG}}$, and red for ${\mathsf{Punt}}$ because we liked Burke’s Twitter thought that fourth-down decision making is like a traffic light. The pink dot (representing the actual play’s yardline and yards to go) lies in a moderately dark green region, indicating a moderate estimated gain in win probability by going for it. Being far from the decision boundary, however, does not imply it the best decision with certainty. Also, it is not obvious how far from the boundary is “far enough.” Hence in Figure 5b we provide an additional chart that illustrates uncertainty in the estimated optimal decision. Here, the color intensity indicates the proportion of bootstrapped models which make the estimated optimal decision. Much of the figure consists of dark colors, specifically dark green in the bottom and middle-left and dark red in the top and middle-right. For those combinations of yardline and yards to go, we are confident in the estimated optimal decision. For example play 2 (the pink dot), ${\mathsf{boot\%}}$ is high (about $90\%$), so we are confident that ${\mathsf{Go}}$ is the best decision. Other combinations of yardline and yards to go feature lighter colors, reflecting high uncertainty. For those game-states, we are not confident in the estimated optimal decision.
Figure 5: Our decision boundary charts for example play 2. The left chart visualizes the estimated optimal decision (color) according to effect size (color intensity) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The pink dot (representing the actual play’s yardline and yards to go) lies squarely in the green region, so the point estimate says ${\mathsf{Go}}$. The right chart is similar except color intensity reflects bootstrap percentage. We are confident that ${\mathsf{Go}}$ is the best decision at this game-state (the pink dot) because it lies in a dark green region. Nonetheless, many other values of yardline and yards to go at this game-state are subject to large uncertainty (lighter color).
We include more examples of our decision procedure in Appendix J. We also include an interactive Shiny App in which you can visualize our fourth-down decision procedure for any game-state on Github.444 https://github.com/snoopryan123/fourth_down/tree/main/3_shiny
4.2 Analytics, have some humility
Since win probability point estimates are subject to substantial uncertainty, we should not simply recommend the decision that maximizes estimated win probability. Due to the random idiosyncrasies in the history of football, which is contained in a relatively small dataset, win probability point estimates can make a decision seem strong even when it isn’t.
In Figure 6 we compare traditional fourth-down recommendations to our recommendations. The strength of a traditional fourth-down decision recommendation is proportional to effect size, the estimated gain in win probability by making that decision ($x$-axis). The strength of our fourth-down recommendation is proportional to the bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). For all fourth-down plays from 2018-2022 in a given effect size bin, we plot the proportion of these decisions we view as confident, lean, and uncertain in Figure 6a. In Figure 6b we plot the frequency of these plays that lie in a given effect size bin.
We are confident in the vast majority of decisions that have an effect size above $4\%$. Most decisions with an effect size under $1\%$ are uncertain, and decisions with an effect size between $1\%$ and $4\%$ are a healthy split between confident, lean, and uncertain. Many plays ($45\%$) have an effect size under $1\%$ and most plays ($87\%$) have an effect size under $4\%$. So, most fourth-down plays feature an effect size that is subject to considerable uncertainty. In particular, we are confident in just $34\%$ of all fourth-down decisions from 2018 to 2022 and a whopping $40\%$ of them are uncertain.
Refer to caption (a) Refer to caption (b)
Figure 6: Left: the proportion of observed fourth-down plays from 2018-2022 that lie in a given decision category (confident, lean, or uncertain according to ${\mathsf{boot\%}}$) as a function of estimated gain in win probability. Right: the frequency of observed fourth-down plays from 2018-2022 that lie in each bucket of estimated gain in win probability.
This analysis reflects substantial overconfidence in win probability point estimates; far fewer fourth-down decisions are as obvious as analysts claim. Although many decisions are estimated to provide solid gains in win probability, many of these perceived edges may actually be due to noise. Thus, we ask football analysts to have some humility: for many game-states there is simply not enough data to be confident in win probability point estimates.
Further, we are confident in about $10\%$ of decisions with effect size under $1\%$ and in about $20\%$ of decisions with effect size between $1\%$ and $2\%$. Although decision confidence is correlated with estimated gain in win probability, they are fundamentally different: we can be confident in plays that provide small edges. Over the course of a season, these edges can accumulate into a large advantage that coaches should take advantage of.
Finally, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, in Appendix I we evaluate coaches only on plays for which we are confident.
Although football analysts have been overconfident in their point estimates for many plays and have been underconfident in certain small edges for other plays, the football analytics community was largely correct that NFL coaches do not go for it enough on fourth down. Across all fourth-down plays from 2018-2022 that we are confident in, the coach made the right decision for $91\%$ of plays where they should have kicked but just $62\%$ of plays where they should have gone for it. Play calling in the NFL is still far too conservative: coaches consistently make wrong decisions, particularly when they should go for it.
5 Discussion
In this study we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a fourth-down recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow that recommendation. Similarly, we believe it is arrogant for a football analyst to expect a coach to follow an uncertain fourth-down recommendation, regardless of the effect size. In today’s age in which sports analysts are fighting a public relations war to be taken seriously by owners and coaches, overconfidence in decision recommendations doesn’t help non-believers buy what we’re selling. Further, the coach is an on-field expert who spends a significant amount of time with his players and has access to information that doesn’t show up in the data (i.e., unobserved confounders). For instance, Eagles coach Nick Sirianni may notice that his usually dominant offensive line is missing a key player today, which does not currently appear in the data. When the estimated optimal fourth-down decision has too much uncertainty, we suggest leaving the decision to the coach’s discretion. Having said that, although football analysts have been overconfident in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
Although our analysis improves the state of the art, it is not without limitations and there are many avenues for future work. We discuss these in detail below and in Appendix K.
Because statistical win probability models produce estimates that are too uncertain at many game-states, we suggest that future work explore probabilistic state-space models to estimate win probability. Probabilistic models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and win probability is calculated by simulating games. The effective sample size of transition probability models is the number of plays because they are fit from independent play-level observations. Some analysts in private industry have created proprietary probabilistic win probability models, which they believe are more accurate than statistical models because they have a larger effective sample size. The tradeoff is that state-space models have higher bias than statistical models. Whereas both classes of models have hidden bias, or bias due to not adjusting for unobserved confounders, just state-space models have model bias as they make simplifying assumptions.
Proprietors of probabilistic models believe that introducing bias in order to reduce variance improves the overall accuracy of the resulting win probability estimator. Nevertheless, these models are subject to their own set of issues, and we believe they aren’t as low-variance as some analysts claim. Properly modeling the distribution of the outcome of a play is nontrivial. In contrast to the simple binary win/loss outcome of statistical win probability models, the outcome variable of a transition probability model is the next game-state, which could include a change in yardline, score, time, or timeouts. This distribution is quite complex: there is a spike at gaining 0 yards for incompletions, a spike for a touchdown, spikes for penalties, and other smooth possibly multimodal distributions for the outcome of run or pass plays, each of which change as a function of team quality and other confounders. Typical transition probability models are riddled with selection bias, as a coach generally chooses play calls that work for his specific players and don’t generalize to an “average” team. Further, uncertainty in these transition probabilities may, after being properly propagated through a state-space model, result in just as much (if not more) uncertainty in estimated win probability than estimates from statistical models. Additionally, one must take great care to carefully encode all the subtle rules of football into her model, and one needs sufficient computing power to simulate enough games to estimate win probability with enough granularity. We look forward to a public facing exploration of probabilistic win probability models in the future.
Acknowledgments
The authors thank the many football analysts who contributed to the development of the fourth down problem. In particular, we thank Brian Burke for providing helpful feedback. The authors acknowledge the High Performance Computing Center (HPCC) at The Wharton School, University of Pennsylvania for providing computational resources that have contributed to the research results reported within this paper.
SUPPLEMENTARY MATERIAL
Appendix A Data details
variable variable description
win/loss 1 if the possession team wins the game, else 0
game seconds remaining num. seconds remaining in the game in $\{3600,...,1\}$
score differential
point differential between the offensive and defensive
team at the start of this play
total score total points scored during this game prior to this play
posteam spread pre-game point spread relative to the possession team
yardline
num. yards from the opponent’s endzone in
$\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
ydstogo
the number of yards the offense has left to gain
before achieving a first down or a touchdown
down a number in $\{1,2,3,4\}$ denoting the down of the play
posteam timeouts remaining num. timeouts in $\{3,2,1,0\}$ the offensive team has
defteam timeouts remaining num. timeouts in $\{3,2,1,0\}$ the defensive team has
receive 2h ko 1 if the possession team receives the $2^{nd}$ half kickoff, else 0
home 1 if the offensive team is at home, else 0
era
categorical variable grouping the year into
[…]
roof
categorical variable grouping the roof of the stadium into
[…]
game id unique identifier of this game
drive id unique identifier of this drive
posteam coach name of the coach of the possession team
Table 3: Game-state variables relevant to estimating win probability that describe the context at the start of a play.
Appendix B Estimating the win probability of a fourth-down decision
In this Section, we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability (denoted $\mathsf{WP_{1}}$) and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We estimate first-down win probability in Section 2.4 and decision transition probabilities in Appendix C.
Win probability of ${\mathsf{Punt}}$. Suppose the offensive team has a fourth down at yardline $y$ and denote the remainder of the game-state by ${\mathbf{x}}$. If the offensive team punts, the opposing team has a first down at the subsequent yardline, which we model as a random variable. Hence the win probability of punting is one minus the opponent’s first-down win probability at the next yardline $y^{\prime}$ and next game-state ${\mathbf{x}}^{\prime}$,
$\sum_{y^{\prime}}(1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime}))\cdot{\mathbb{P}}(\text{yardline after punting is $y^{\prime}$}|{\mathbf{x}}).$ (11)
The next game-state ${\mathbf{x}}^{\prime}$ flips the current game-state variables in ${\mathbf{x}}$ that are relative to the team with possession (e.g., score differential, team quality metrics, timeouts remaining, etc.) and doesn’t alter the other variables (e.g., time remaining). We re-write this expression in terms of the expectation over the outcome of the punt,
${\mathbb{E}}_{\text{punt}}[1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime})|{\mathbf{x}}].$ (12)
First-down win probability is mostly linear in yardline at most game-states, so for simplicity we instead compute win probability at the expected next yardline after punting,
$1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{punt}}[y^{\prime}|{\mathbf{x}}],{\mathbf{x}}^{\prime}).$ (13)
We model the expected next yardline after punting as a function of yardline and punter quality in Appendix C.1.
Win probability of ${\mathsf{FG}}$. We decompose field goal win probability in terms of field goal success probability and first-down win probability on the subsequent play,
[…]
We model the probability of a successful field goal as a function of yardline and kicker quality in Appendix C.2. If the kicking team misses the field goal and the spot of the kick was within the 20 yardline, the opposing team has a first down at yardline $80$. If the kicking team misses the field goal and the spot of the kick was beyond the 20 yardline, the opposing team has a first down at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
[…]
If the kicking team makes the field goal, it scores three points and the opposing team has a first down after a kickoff. Denoting the score differential relative to the kicking team by $s$, we have
${\mathsf{WP}}(\text{make FG})=1-{\mathbb{E}}_{\text{kickoff}}[\mathsf{WP_{1}}(\text{yardline }y^{\prime},s^{\prime}=-s-3,{\mathbf{x}}^{\prime})].$ (16)
As we did in estimating ${\mathsf{Punt}}$ win probability, for simplicity we instead compute win probability at the expected next yardline after a kickoff,
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{kickoff}}[y^{\prime}],\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (17)
The vast majority of kickoffs end in a touchback (yardline $75$), so for simplicity we instead compute
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }75,\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (18)
Win probability of ${\mathsf{Go}}$. Suppose the offensive team has a fourth down and $z$ yards-to-go at yardline $y$. If the offensive team goes for it and gains $\Delta\geq z$ yards, then in the next play it either has a first down at yardline $y-\Delta$ or a touchdown. Conversely, if the offensive team goes for it and gains $\Delta<z$ yards, then in the next play the opposing team has a first down at yardline $100-(y-\Delta)$. Hence the expected value of going for it on fourth down is
$\sum_{\Delta\geq z}\mathsf{WP_{1}}(\text{yardline }y-\Delta)\cdot{\mathbb{P}}(\text{gain $\Delta$ yards})+\sum_{\Delta<z}(1-\mathsf{WP_{1}}(\text{yardline }100-(y-\Delta)))\cdot{\mathbb{P}}(\text{gain $\Delta$ yards)}.$) . (19)
Here, the first-down win probability and conversion success probability are implicitly functions of game-state. As the probability distribution of the yards gained on a conversion attempt is complex, we employ a standard simplification from Burke, 2009b . If the offensive team converts on fourth down and $z$ yards-to-go at yardline $y$, we assume they gain $z$ yards on that play, yielding a first down at yardline $y-z$ or a touchdown. If the offensive team fails to convert, we assume they turn the ball over at the initial yardline of the play, leaving the opposing team with a first down at yardline $100-y$. Thus, the win probability of going for it is
$\begin{split}\mathsf{WP_{1}}(\text{yardline }y-z,{\mathbf{x}}^{\prime})\cdot{\mathbb{P}}(\text{convert}|{\mathbf{x}})+(1-\mathsf{WP_{1}}(\text{yardline }100-y,{\mathbf{x}}^{\prime}))\cdot(1-{\mathbb{P}}(\text{convert}|{\mathbf{x}})),\end{split}$ (20)
where ${\mathbf{x}}$ is the game-state on fourth down and ${\mathbf{x}}^{\prime}$ is the game-state on the subsequent play (as before, ${\mathbf{x}}^{\prime}$ flips the game-state variables which are relative to the team with possession (e.g., score differential, team quality metrics, timeouts remaining) and doesn’t alter the other variables (e.g., time remaining)). We model conversion probability as a function of yardline, yards-to-go, and offensive and defensive quality in Appendix C.3.
Appendix C Estimating fourth-down decision transition probabilities
In this Section we estimate fourth-down decision transition probability models. In Section C.1 we model the expected next yardline after punting, in Section C.2 we model field goal success probability, and in Section C.3 we model conversion success probability.
C.1 Modeling punt outcome
In this Section we model the outcome of a punt. Specifically, we model the expected next yardline after a punt (from the perspective of the receiving team) as a function of yardline and punter quality (from the perspective of the punting team).
Punter quality. We define a punter’s quality by a weighted sum of his punt yards over expected over all his previous punts in his career. To begin, we fit a simple expected next yardline after punting model ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ on a held-out dataset of all punts from 1999 to 2005 to avoid data bleed. We fit ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ using linear regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the punt yards over expected (${\mathsf{PYOE}}$) of the $n^{th}$ punt by
${\mathsf{PYOE}}_{n}:=\text{actual yardline after the }n^{th}\text{ punt}-{{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}(\text{yardline prior to the }n^{th}\text{punt}).$ (21)
Now we define punter quality using Rams’ punter Johnny Hekker for concreteness. Index all of Hekker’s punts from 2006 to 2021 by $n$. We define Hekker’s punter quality prior to punt $n$ by a weighted sum of the punt yards over expected from his previous kicks,
${\mathsf{pq}}_{n}:=\alpha\cdot{\mathsf{pq}}_{n-1}+{\mathsf{PYOE}}_{n-1},$ (22)
where ${\mathsf{pq}}_{0}:=0$ and ${\mathsf{PYOE}}_{0}:=0$. We use an exponential decay weight $\alpha=0.995$ to upweight more recent punts. Finally, we standardize the punter quality column to have mean zero and standard deviation $1/2$. In Figure 7a we plot the career mean punter quality of each punter with over 250 punt attempts from 2006 to 2021. As expected, Johnny Hekker has the highest punter quality.
Punt outcome model. We use linear regression to model the expected next yardline after a punt as a function of yardline and punter quality (${\mathsf{pq}}$),
$\begin{split}{\mathbb{E}}_{\text{punt}}[\text{next yardline}]&=\ \vec{\alpha}\cdot\text{spline}\big{(}\text{yardline},\ \text{df}=4\big{)}+\ \beta_{1}\cdot{\mathsf{pq}}+\beta_{2}\cdot{\mathsf{pq}}\cdot\text{yardline}.\end{split}$ (23)
The model is trained on a dataset of $36,493$ punts from 2006 to 2021, all beyond the 30 yardline. In Figure 7b we plot the expected next yardline after a punt according to our model as a function of yardline for various punter quality values.
Figure 7: On the left, the career mean punter quality ($x$-axis) of each punter ($y$-axis) with over 250 punts from 2006 to 2021. On the right, the expected next yardline after a punt ($y$-axis) according to our model as a function of yardline ($x$-axis) for various punter quality values (color).
C.2 Modeling field goal success probability
In this Section we model field goal success probability as a function of yardline and kicker quality. It is important to adjust for kicker quality to avoid selection bias, as good kickers attempt more field goals from long distance than bad kickers. Our kicker quality adjustment, detailed below, is similar to our punter quality adjustment from the previous Section.
Kicker quality. We define a kicker’s quality by a weighted sum of his field goal probability added over all his previous kicks in his career. To begin, we fit a simple field goal probability model ${P_{\mathsf{FG}}^{(0)}}$ on a held-out dataset of all field goals from 1999 to 2005 to avoid data bleed. We fit ${P_{\mathsf{FG}}^{(0)}}$ using logistic regression as a function of yardline (specifically, a cubic polynomial spline with five degrees of freedom on the yardline). Then, we define the field goal probability added (${\mathsf{FGPA}}$) of the $n^{th}$ field goal by
${\mathsf{FGPA}}_{n}:=\mathbb{I}\left(n^{th}\text{ field goal is made}\right)-{P_{\mathsf{FG}}^{(0)}}(\text{yardline of the }n^{th}\text{ field goal}).$ (24)
Now we define kicker quality using the Ravens’ kicker Justin Tucker for concreteness. Index all of Tucker’s field goals from 2006 to 2021 by $n$. We define Tucker’s kicker quality prior to field goal $n$ by a weighted sum of the field goal probability added in his previous kicks,
${\mathsf{kq}}_{n}:=\alpha\cdot{\mathsf{kq}}_{n-1}+{\mathsf{FGPA}}_{n-1},$ (25)
where ${\mathsf{kq}}_{0}:=0$ and ${\mathsf{FGPA}}_{0}:=0$. As before, we use an exponential decay weight $\alpha=0.995$ to upweight more recent kicks. Finally, we standardize the kicker quality column to have mean zero and standard deviation $1/2$. In Figure 8a we plot the career mean kicker quality of each kicker with over 100 field goal attempts from 2006 to 2021. As expected, Justin Tucker has by far the highest kicker quality.
Field goal success probability model. We use logistic regression to model the probability that a kicker makes a field goal as a function of yardline and kicker quality (${\mathsf{kq}}$). Formally, our best field goal probability model is
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}(\text{make }{\mathsf{FG}})}{1-{\mathbb{P}}(\text{make }{\mathsf{FG}})}\bigg{)}=\ \vec{\alpha}\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=5\big{)}+\beta\cdot{\mathsf{kq}}.\end{split}$ (26)
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. In Figure 8b we plot the probability of making a field goal according to our model as a function of yardline for various kicker quality values.
Figure 8: On the left, the career mean kicker quality ($x$-axis) of each kicker ($y$-axis) with over 100 field goal attempts from 2006 to 2021. On the right, the probability of making a field goal ($y$-axis) according to our model as a function of yardline ($x$-axis) for various kicker quality values (color).
C.3 Modeling conversion probability
In this Section we model conversion probability as a function of yards to go and team quality. First, we detail our team quality measures, which produce a more predictive conversion probability model than point spread. Then, we detail our conversion probability model.
Team quality. We create measures of offensive and defensive quality using the result of each previous play. A common and good way of quantifying play success is expected points added (${\mathsf{EPA}}$) (Yurko et al.,, 2018). The ${\mathsf{EPA}}$ of play $n$ is the difference in expected points (${\mathsf{EP}}$) between the end and the beginning of the play,
${\mathsf{EPA}}_{n}={\mathsf{EP}}_{n}-{\mathsf{EP}}_{n-1}.$ (27)
To avoid data bleed, we remove all plays from 1999 to 2005 from our play-by-play dataset and fit an expected points model ${\mathsf{EP^{(0)}}}$ to this held-out data.
The simple ${\mathsf{EP^{(0)}}}$ model which we fit on held-out data from 1999-2005 in order to craft our team quality metrics is a linear regression in which the outcome variable is the observed net points of the next score, a real number in $\{7,3,2,0,-2,-3,-7\}$. We model expected net points as a linear function of yardline, log yards to go, and one combined indicator for third and fourth down.
Figure 9: The career mean quarterback quality ($x$-axis) of each quarterback ($y$-axis) with over 1750 attempts and whose rookie season came after 2006.
Then we use ${\mathsf{EPA^{(0)}}}$ (${\mathsf{EPA}}$ derived from ${\mathsf{EP^{(0)}}}$) to craft our team and player quality metrics. For concreteness, consider deriving the quarterback quality of Patrick Mahomes. Index all the plays from 2006 to 2021 in which Mahomes passes or runs the ball by $n$. We define Mahomes’ quarterback quality prior to play $n$ by a weighted sum of the ${\mathsf{EPA^{(0)}}}$ of his previous plays,
$q_{n}:=\gamma_{n}\cdot\alpha\cdot q_{n-1}+{\mathsf{EPA^{(0)}}}_{n-1},$ (28)
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays. We set $\alpha=0.995$ for each of our team and player quality metrics. For instance, a play which occurred 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occurred 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 4. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noisier from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 9 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 4 in a similar fashion, via Formula (28).
variable $\gamma$
quarterback quality $3/4$
non-quarterback offensive quality $1/2$
defensive quality against the pass $1/3$
defensive quality against the run $1/3$
Table 4: Summary of team quality metrics computed via Formula (28) and their start-of-season shrinkage parameters $\gamma$. Each of these metrics apply to both the offensive and defensive teams in a play, yielding eight total metrics.
Due to a small sample size of fourth-down conversion attempts, some existing models use third down as a proxy for fourth down, as they are also high-pressure situations in which the offensive team attempts to gain a first down on that play (Romer,, 2006). There may be, however, a fundamental difference in conversion probability between third and fourth-down plays, perhaps due to psychological reasons or a different distribution of play calls on fourth down. Therefore, in our model selection process, we test models on a random 50% of fourth-down plays, and we train some models on a dataset consisting entirely of fourth-down plays and other models on a dataset consisting of third and fourth-down plays. We find that the parameters of our best conversion probability model borrow strength from third-down plays.
Our best logistic regression model adjusts for yards to go, down (third vs. fourth down), and our offensive and defensive quality metrics: quarterback quality of the offensive team (${\mathsf{qbqot}}$), non-quarterback offensive quality of the offensive team (${\mathsf{oqrot}}$), defensive quality of the defensive team against the pass (${\mathsf{dqdtap}}$), and defensive quality of the defensive team against the run (${\mathsf{dqdtar}}$). Formally, our best conversion probability model is
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (29)
In Figure 10 we visualize conversion probability as a function of yards to go. We see a large spike in conversion probability with one yard to go. In Figure 11 we plot conversion probability as a function of yards to go for various values of team quality. We find that quarterback quality significantly impacts conversion probability and the other team quality measures have a small impact.
Figure 10: Conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis), assuming average team quality.
Figure 11: Conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis) for various values of team quality (color). Quarterback quality has by far the largest impact on conversion probability, visualized by a larger spread amongst the colored lines.
A more elaborate conversion probability model may adjust for yardline. In particular, it is plausible that it is more difficult to convert near each endzone, where space is more constricted. Additionally, a more fine-grained model would be continuous in yards to go rather than treating it as an integer. Of course, our model can only be as good as available data, which treats yards to go as an integer. But, anecdotally, fourth down and inches has a significantly higher conversion probability than fourth down and one yard to go. Finally, conversion probability depends on the offensive and defensive play call. On this view, in practice it may be better for teams to input their own conversion probability estimates as they are more informed on their play calling tendencies.
Appendix D First-down win probability model details
In our fourth-down decision procedure, we use a win probability ${\mathsf{XGBoost}}$ model fit from first-down plays. This model predicts binary win/loss as a function of score differential, game seconds remaining, point spread, yardline, receive $2^{nd}$ half kickoff indicator, offensive team’s number of timeouts remaining, defensive team’s number of timeouts remaining, total score, and
$\texttt{scoreTimeRatio}=\frac{\texttt{score\_differential}}{0.01+\texttt{game\_seconds\_remaining}}.$ (30)
We include monotonic increasing constraints for score differential, scoreTimeRatio, and offensive timeouts remaining and monotonic decreasing constraints for point spread, yardline, and defensive timeouts remaining.
Now we compare the out-of-sample predictive performance of various ${\mathsf{WP}}$ models. Our full dataset consists of all football plays from 2006 to 2021. The dataset is clustered into games, as plays from each game share the same winning team. To keep the clustered nature of our dataset intact and to avoid data bleed, we split our dataset in half by randomly sampling 50% of all games. The first-down plays from the first 50% of these games form the hold-out test set. We test on first-down plays because fourth-down decision making relies on first-down win probability. The plays from the other 50% of these games form the training set. To tune ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We visualize the results of our model comparison in Table 5.
model
name
model
type
plays
in dataset
out-of-sample
${\mathsf{logloss}}$
Ours ${\mathsf{XGBoost}}$ first downs 0.440
Lock and Nettleton, (2014) ${\mathsf{RF}}$ all downs 0.447
Baldwin, (2021) ${\mathsf{XGBoost}}$ all downs 0.476
Table 5: Predictive performance of various ${\mathsf{WP}}$ models.
Appendix E Simulation study details
Generating plays. Formally, the outcome of the $n^{th}$ play of the $g^{th}$ game is
$\xi_{gn}\overset{iid}{\sim}\pm 1.$ (31)
The game starts at midfield, $X_{g0}=L/2$, and the game begins tied, $S_{g0}=0$. The field position at the start of play $n$ is
$X_{g,n+1}:=\begin{cases}X_{gn}+\xi_{gn}&\text{ if }0<X_{gn}+\xi_{gn}<L\text{ (not a TD)}\\ L/2&\text{ else,}\end{cases}$ (32)
and the score differential at the start of play $n$ is
$S_{g,n+1}:=\begin{cases}S_{gn}+1&\text{ if }X_{gn}+\xi_{gn}=0\text{ (TD)}\\ S_{gn}-1&\text{ if }X_{gn}+\xi_{gn}=L\text{ (opp. TD)}\\ S_{gn}&\text{ else.}\end{cases}$ (33)
The response column win is
$y_{gn}\equiv y_{g,N+1}:=\begin{cases}1&\text{ if }S_{g,N+1}>0\\ 0&\text{ if }S_{g,N+1}<0\\ \text{Bernoulli}(1/2)&\text{ else (overtime). }\end{cases}$ (34)
As in our dataset of real football plays, this response column is highly correlated – plays from the same game share the same draw of the winner of the game.
Generating observational data. We create a dataset of plays from $G$ games. Each game consists of $N$ plays, and the field consists of $L$ yardlines. The results from each game yield a simulated dataset
${\mathscr{D}}=\{(n,X_{gn},S_{gn},y_{gn}):n=1,...,N\text{ and }g=1,...,G\}.$ (35)
True win probability. The true win probability
${\mathsf{WP}}(n,x,s):={\mathbb{P}}(S_{g,N+1}>0|X_{gn}=x,S_{gn}=s)$ (36)
of our simplified version of football is computed explicitly using dynamic programming,
${\mathsf{WP}}(N+1,x,s)=\begin{cases}1&\text{if }s>0\\ 1/2&\text{if }s=0\\ 0&\text{if }s<0,\\ \end{cases}$ (37)
and
${\mathsf{WP}}(n-1,x,s)=\begin{cases}\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s+1)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{if }x=1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s-1)&\text{if }x=L-1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{else.}\end{cases}$ (38)
Appendix F Fitting our fourth-down decision procedure
In this Section we describe in detail our fourth-down decision procedure, which incorporates uncertainty quantification.
We begin by fitting the traditional fourth-down decision procedure. We fit our first-down win probability model, an ${\mathsf{XGBoost}}$ model fit from first-down plays, to our dataset (recall Section 2.4 and Appendix D). We also fit our decision transition models from our dataset (recall the punt outcome model, field goal success probability model, and conversion probability model, described in Appendix C). We then knit these models together to estimate fourth-down win probability for each decision ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ (described in Section 2 and Appendix B). The estimated optimal decision is the one that maximizes win probability.
We use bootstrapping to quantify uncertainty in our fourth-down decision recommendations. We generate $B=100$ bootstrapped datasets using a randomized cluster bootstrap as follows (we justify our choice of $B=100$ in Appendix G). Let $G$ denote the number of games in our observed play-by-play dataset. To generate a bootstrapped dataset we re-sample $G\cdot\phi$ games uniformly with replacement. We use $\phi=0.5$ as justified by the simulation study from Section 3.1 in order to achieve adequate marginal coverage. Within each game $g$, which contains $N_{g}$ drives, we re-sample $N_{g}$ drives uniformly with replacement. We re-sample drives with replacement, rather than plays with replacement as we did in the simulation study, since it provides a more realistic re-sampling of real football data.
From each of the $B$ bootstrapped datasets we fit a first-down win probability model. We knit each bootstrapped first-down win probability model together with the decision transition models fit from observed data to form $B$ bootstrapped fourth-down decision models. For each game-state, from the $B$ bagged decisions we calculate bootstrap percentage and form a $90\%$ confidence interval on the estimated gain in win probability.
We just bootstrap the first-down win probability model and not the decision transition models because the former model is subject to much higher uncertainty. Decision transition probability models are fit from i.i.d. play-level outcomes, whereas first-down win probability models are fit from clustered, highly correlated game-level outcomes. On this view, as discussed in Section 2.3, estimating first-down win probability is the most difficult component of fourth-down decision making and is the source of the majority of uncertainty in fourth-down win probability estimates. Hence in this paper we focus on propagating uncertainty from first-down win probability models into the fourth-down decision procedure.
Appendix G Tuning the number of bootstrapped datasets
As discussed in Section 3.2, we use bootstrap percentage to incorporate uncertainty quantification into fourth-down decision making because it quantifies uncertainty in the decision itself. Although the strength of a decision is proportional to the ${\mathsf{boot\%}}$, we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a decision recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
Since we ultimately bucket decision recommendations into three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
Across $M=100$ draws of $B$ bootstrapped decision models in each, we categorize each observed fourth-down play from 2018 to 2022 into one of the three bins. Let $i$ index all the observed fourth-down plays from 2018 to 2022, let $m\in\{1,...,M\}$ index the draw of the bootstrap, and in each draw of the bootstrap we generate $B$ fourth-down decision models. Given ${\mathsf{boot\%}}_{im}^{(B)}$, the ${\mathsf{boot\%}}$ of the estimated optimal decision for play $i$ in draw $m$ of the bootstrap, we calculate
$\begin{cases}p_{i}^{\text{confident, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[90\%,100\%]\right),\\ p_{i}^{\text{lean, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[75\%,90\%)\right),\\ p_{i}^{\text{uncertain, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[0\%,75\%)\right),\\ p_{i}^{(B)}=\max\{p_{i}^{\text{confident, }(B)},\ p_{i}^{\text{lean, }(B)},\ p_{i}^{\text{uncertain, }(B)}\}.\end{cases}$ (39)
For each category, $p_{i}^{\text{category, }(B)}$ is the proportion of the $M$ bootstrap draws that play $i$ is put in that category. For our procedure with $B$ bootstrapped models to be stable, we want $p_{i}^{\text{category, }(B)}$ to be close to 1 for each category, which means that a play’s categorization is not dependent on the randomness inherent in generating $B$ bootstrapped models. Hence, we want $p_{i}^{(B)}$, the maximum $p_{i}^{\text{category, }(B)}$ across the three categories, to be close to 1. Hence, we want the mean across all plays $\overline{p}^{(B)}=\frac{1}{n}\sum_{i=1}^{n}p_{i}^{(B)}$ to be close to 1. In Figure 12 we see that for $B=100$, $\overline{p}^{(B)}=0.9$ and the vast majority of plays $i$ have $p_{i}^{(B)}=1$. Some plays have $p_{i}^{(B)}$ lower than 1 since they lie near the border of two categories. We believe $\overline{p}^{(B)}=0.9$ is sufficiently large and hence use $B=100$ in our fourth-down decision procedure.
Figure 12: This is a histogram of $\{p_{i}^{(B)}\}_{i}$. The blue line is the mean. Since most of the value equals 1 and the mean is high enough, $B=100$ is large enough.
Appendix H Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is an independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 13 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.
Figure 13: Visualizing our model of the typical coach’s fourth-down decision as a function of yardline and yards to go, for various values of time remaining and score differential. Green, yellow, and red indicate that ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ is the most likely decision, respectively. The color intensity reflects the likelihood that a coach makes that decision.
In Figure 14 we visualize the variable importance (via gain) of our ${\mathsf{XGBoost}}$ model. Interestingly, point spread has an extremely small impact on coaches’ fourth-down decisions. Perhaps this is because coaches don’t like to admit when their teams are underdogs as some sort of psychological leadership tool. We find, however, that point spread should impact fourth-down decision making. For instance, in certain game-states, it is advantageous for the favorites to be more aggressive (e.g., late in close games).
Figure 14: Variable importance (gain) for the typical coach’s decision probability model.
Appendix I Evaluating coaches’ decision making
A football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. On this view, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 15 we rank coaches from 2018 to 2022 by the proportion of correctly made decisions that we are confident in. The top of this list has coaches like Kevin O’Connell and John Harbaugh from analytics bent organizations but also coaches that are considered more traditional on fourth down such as Ron Rivera, Sean McVay, and Mike Vrabel. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
Figure 15: Ranking coaches by the proportion of fourth-down decisions they made correctly from 2018-2022 among fourth-down plays we are confident in.
Appendix J Additional example plays
Example play 3. In Figure 16 we visualize our decision procedure for a fourth-down play in which the Bears had the ball against the Jets in Week 12 of 2022. ${\mathsf{FG}}$ provides an edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.3\%$ ${\mathsf{WP}}$). But our $90\%$ confidence interval of the estimated gain in win probability by attempting a field goal is $[-3.2\%,3.9\%]$, indicating that ${\mathsf{FG}}$ could either be a great or a terrible decision. Also, more than half of our bootstrapped models say ${\mathsf{Go}}$ is better. In other words, we do not have enough data to be confident in our win probability point estimates, and we don’t know the optimal fourth-down decision at this game-state. Further, in the bottom right plot, notice how most of the colors are light. This indicates that the optimal decision is uncertain at most other combinations of yardline and yards to go at this game-state.
Refer to captionRefer to caption
Figure 16: Our decision charts for example play 3. The bootstrapped models are split amongst ${\mathsf{Go}}$ and ${\mathsf{FG}}$ (${\mathsf{boot\%}}$, in the orange column in the top figure, is near $50\%$-$50\%$), so we don’t have enough data to be confident in the fourth-down recommendation provided by the point estimate (blue column in the top figure). Even though the pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a yellow region (so the point estimate says ${\mathsf{FG}}$), the pink dot in the lower right figure is in a region of light color intensity (reflecting low ${\mathsf{boot\%}}$).
Example play 4. In Figure 17 we visualize our decision procedure for a fourth-down play in which the Ravens had the ball against the Jets in Week 1 of 2022. ${\mathsf{Punt}}$ provides a slight edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.2\%$ ${\mathsf{WP}}$). But $97\%$ of the bootstrapped models say ${\mathsf{Punt}}$ is better and our $90\%$ confidence interval of the estimated gain in win probability by punting is $[0.5\%,3.5\%]$, which is strictly positive. So, even if the edge is small, we are confident in this edge and recommend that the Ravens should ${\mathsf{Punt}}$. Further, in the bottom right plot, notice how most of the colors are dark outside of a large white boundary region. This indicates that we have higher certainty in our estimated optimal decision at this game-state.
Refer to captionRefer to caption
Figure 17: Our decision charts for example play 4. Nearly all the bootstrapped models say ${\mathsf{Punt}}$ is superior (${\mathsf{boot\%}}$, in the orange column in the top figure, is $97\%$), so we are confident that ${\mathsf{Punt}}$ has higher win probability than the other decisions. The pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a red region (so the point estimate says ${\mathsf{Punt}}$) and the pink dot in the lower right figure is in a region of dark color intensity (reflecting high ${\mathsf{boot\%}}$).
Example play 5. In Figure 18 we visualize our decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+3.9\%$ ${\mathsf{WP}}$). Further, $92\%$ of the bootstrapped models say ${\mathsf{Go}}$ is better and our $90\%$ confidence interval of the estimated gain in win probability by going for it is $[-0.1\%,4.4\%]$, which is nearly positive. Thus, we are confident in this edge, and we strongly recommend that the Raiders should ${\mathsf{Go}}$555 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98-yard drive to win the game.
Refer to captionRefer to caption
Figure 18: Our decision charts for example play 5. Nearly all the bootstrapped models say ${\mathsf{Go}}$ is superior (${\mathsf{boot\%}}$, in the orange column in the top figure, is $92\%$), so we are confident that ${\mathsf{Go}}$ has higher win probability than the other decisions. The pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a green region (so the point estimate says ${\mathsf{Go}}$) and the pink dot in the lower right figure is in a region of dark color intensity (reflecting high ${\mathsf{boot\%}}$).
Appendix K Limitations and future work
Although our analysis improves the state of the art, it is not without limitations and there are many avenues for future work. To begin, even though the randomized cluster bootstrap produces substantially wide confidence intervals for win probability estimates, it underestimates uncertainty because it quantifies sampling uncertainty but not model uncertainty. The former is uncertainty in our point estimates resulting from fitting a model on a finite dataset (“variance”) and the latter is uncertainty caused by our model being wrong or biased (“bias”). In our simulation study from Section 3.1, there is no model uncertainty because we know true win probability is indeed a function of time, score differential, and field position. Win probability in real football, however, is highly likely a function of unobserved confounders. For example, how well Tom Brady slept the previous night could affect his team’s win probability. A more elaborate uncertainty quantification would capture model uncertainty.
In this work we used bootstrapping to drive uncertainty quantification. The way we tuned the bootstrap to achieve adequate marginal coverage in Section 3.1 is reminiscent of conformal inference. Conformal prediction intervals provide guarantees regarding the marginal coverage by ensuring every observation gets a holdout version of the interval (Lei et al.,, 2018). That the bootstrap methods with $\phi=1$ don’t achieve adequate coverage could be due to the time series dependence structure of football data. Recent work on conformal approaches for time series may help mitigate this issue (Angelopoulos et al.,, 2023). Nonetheless, valid distribution-free inference may not be possible in a binary outcome setting (Barber,, 2020). For instance, no confidence or prediction interval (e.g., $[0.63,0.71]$) will cover a 00 or $1$ outcome value.
The way we tuned the bootstrap is also reminiscent of random forest prediction intervals (Zhang et al.,, 2020). Random forest prediction intervals are based on the empirical distribution of out-of-bag prediction errors. These intervals can be obtained as a by-product of a single random forest and under regularity conditions have asymptotically correct coverage rates.
Our analysis captures uncertainty in first-down win probability but not in decision transition probabilities (conversion probability, field goal success probability, and expected next yardline after punting). These models are fit from thousands of play-level i.i.d. observations, and so are subject to little sampling uncertainty, but are subject to nontrivial model uncertainty, as they make simplifying assumptions. Conversion probability in particular is a delicate concept, as it depends on the offensive play call, the defensive play call, and the individual characteristics of each of the players on the field. A more fine-grained analysis would account for this additional uncertainty.
Finally, in this work we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If teams follow our recommendations, their behavior will change and win probability will change accordingly. Statistical win probability models that learn from the game outcomes defining the recent history of football do not account for this distribution shift. State-space models, on the other hand, can account for these changes by altering the probability that a team goes for in on fourth down as a function of game-state. A more elaborate analysis would account for this distribution shift.